What is the integration of 7^xe^x dx

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
\\ \\y=(7e)^x \\$This is same as integration of $A^x \\\int y=\frac{A^x}{\ln \left(A\right)}+C \\
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty